OpenStudy (anonymous):

show me how to solve for x when 1/(x+1)^2 = [x/(x+1) - 2]/(x-1)

OpenStudy (anonymous):

bunch of algebra, i will try to write it

OpenStudy (anonymous):

ik thanks

OpenStudy (anonymous):

\[\frac{x}{x+1}-2=\frac{-x-2}{x+1}\] is a start

OpenStudy (anonymous):

so right hand sides is \[\frac{-x-2}{(x+1)(x-1)}\]

OpenStudy (anonymous):

wait how did you start out like that

OpenStudy (anonymous):

\[\frac{1}{(x+1)^2}=\frac{-x-2}{(x+1)(x-1)}\] \[\frac{1}{x+1}=\frac{-x-2}{x-1}\] \[x-1=(x+1)(-x-2)\] \[x-1=-x^2-3x-2\] \[x^2+3x+1=0\]

OpenStudy (anonymous):

\[\frac{x}{x+1}-2=\frac{x}{x+1}-\frac{2(x+1)}{x+1}\] etc

OpenStudy (anonymous):

still don't get thte first step talk about moronic what the flutter